OpenStudy (anonymous):

16. A ferris wheel is 20 meters in diameter and boarded from a platform that is 4 meters above the ground. The six o'clock position on the ferris wheel is level with the loadin gplatform. The weheel completes one full revolution every 2 minutes. at t=0 you are in the 12 o'clock position. you then make two complete revolution needed to return to the boarding platform. graph f=h(t), first determine an appropriate interval for t, with t greater than or equal to 0 Label the period, the amplitutde, and the midline of each graph,

OpenStudy (anonymous):

i am confused...if you start at the 12 o'clock position...then you need 2 and a half revolutions to get back to the platform, not 2? period = 2 amplitude = 10 midline = 4 + 10 = 14 t interval [0, 5]

OpenStudy (anonymous):

why time interval [0,5] not 4? it indicates 2 complete revolutions not 2.5

OpenStudy (anonymous):

i did get this. I'm having trouble finding out the formula for this.

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

beehive, if you start at the 12 o'clock position, and you want to end up in the 6'oclock position (which is where the boarding platform is), then two full revolutions will just lead you back to the 12 o'clock position

OpenStudy (anonymous):

true, i didn't look at it this way.

OpenStudy (anonymous):

sketch a sinusoidal curve that goes from 4 to 24 with a period of 2

OpenStudy (anonymous):

lol, jamesm are you beehive's teacher? lol

OpenStudy (anonymous):

that be cool. But all i need right now is to understand what i need to do in this chapter and finish homework. I got 2 more chapters of similar stuff to do. and i don't know anyting.

OpenStudy (anonymous):

jamesm. I know how the graph looks like. Sin function and so one. I don't know how to make it into a function, give the above facts about it.

OpenStudy (anonymous):

well if you know what it looks like...and you know what the basic graphs y = sin(x) and y = cos(x) looks like...what has been done to those basic graphs to make it look like the ferris wheel graph?

OpenStudy (anonymous):

jamesm go http://www.twiddla.com/solved